### v2

@@ Abstract
for arbitrary values of $N$, $n$, $\epsilon$ and $p^{\rm th}$. Naturally, an interplay exists between $\epsilon$ and $q^{\rm th}$: fulfilling Eq. (1) with a very small $\epsilon$ comes at the price of tolerating larger values of $q^{\rm th}$. In the context of QKD —in which $\hat{q}$ is often referred to as the phase-error rate (PHER)— this translates into a direct penalty on the provably secure key length. Notably, the above formulation of the problem shows up exactly in ideal implementations of the BB84 [4] or the BBM92 [7, 8] protocols, and we adhere to the latter in the main text for didactic purposes. The detailed protocol description and the secret key length formula can be found in Supplemental Material I. On the other hand, the popular decoy-state BB84 protocol [9, 10, 11, 12], which is implemented in most of today’s commercial QKD setups, is addressed in Supplemental Material II.
@@
Crucially, our relaxation stems from a rational approximation of the Kullback-Leibler divergence, which is based on a logarithmic inequality [24] carefully selected to keep the CIs tight and handy at the same time. Indeed, higher-order rational approximations of the logarithm lead to much more convoluted formulas for the divergence. The reader is referred to Supplemental Material III for the relaxed Chernoff bound, and the corresponding CIs are stated in Proposition 1 and Corollary 1 below.
@@
A proof of this claim is given in Supplemental Material III. Importantly, since the claim is solely derived from the additive Chernoff bound, it automatically extends to a HG random variable [17, 21]. The upper extension goes as follows.
@@
On another note, the reader is referred to Supplemental Material II for the failure probability estimation of a decoy-state QKD protocol, which relies not only on Proposition 2 but also on Proposition 1.
@@
On another note, in Supplemental Material IV we include optimal one-sided CIs for the average of independent Bernoulli parameters [30, 31], potentially replacing $\Gamma^{\pm}_{n,\epsilon}(\hat{p})$ (Proposition 1) in the failure probability estimation of decoy-state QKD schemes. These CIs are related to the well-known Clopper-Pearson CIs of the binomial distribution, which in fact have been used in the context of QKD security before (see e.g. [32, 33]).
QKD simulations. Next, we show the impact that the proposed tools have in the finite-key performance of QKD. Let us consider the ideal BBM92 protocol first. Consistently with [15, 8], we evaluate a simplified scenario in which Alice and Bob share a $2N$-partite quantum state as a protocol input. Then, they jointly select between two binary and complementary quantum measurements uniformly at random on a round-by-round basis, in so each of them extracting a raw data block of $N$ bits. These blocks are randomly splitted into a fixed-length test (sifted-key) string of $n$ ($N-n$) bits. That is to say, both the key and the test string are randomly sampled from the two bases. In the simulations, we set the tolerated quantum bit error rate (QBER) to $p^{\rm th}=4.55\%$ for ease of comparison with [15], where this choice is motivated by the results of the Micius experiment [34]. Also, an error correction (EC) leakage of $\lambda_{\rm EC}=1.19(N-n)h(p^{\rm th})$ bits is assumed, $h(x)$ denoting the binary entropy of $x$. The correctness parameter, $\epsilon_{\rm cor}$, and the privacy amplification error, $\epsilon_{\rm PA}$, are both set to $10^{-8}$ for illustration purposes, and the failure probability bound is set to $\epsilon_{\rm PE}=4\times{}10^{-16}$. Altogether, these settings lead to an overall secrecy parameter of $\epsilon_{\rm sec}\approx{}5\times{}10^{-8}$ (see Supplemental Material I).
@@
Let us now assess the more practical decoy-state BB84 protocol. We consider a scheme with three common intensities per basis, $\mu$ (signal), $\nu$ (decoy) and $\omega$ (vacuum), and employ the $Z$ ($X$) basis for key extraction (PE) with a fixed-size data block of $N_{\rm Z}$ ($N_{\rm X}$) bits drawn at random from the corresponding data pool. Again, in the simulations we contemplate blocks up to $N=N_{\rm Z}+N_{\rm X}=10^{6}$ bits, and superpose an inset of the minimum block-size regime. Naturally, we set $N_{\rm Z}=q_{\rm Z}^{2}/(q_{\rm Z}^{2}+q_{\rm X}^{2})N$, where $q_{\rm Z}$ ($q_{\rm X}$) denotes the $Z$ ($X$) basis probability, common to both parties. Also, while we fix $\omega=10^{-4}$ to account for the finite extinction ratio of the intensity modulator [35], the settings $\mu$ and $\nu$, together with their respective probabilities, $p_{\mu}$ and $p_{\nu}$, and the test basis probability, $q_{\rm X}$, are tuned to optimize the key length per block, $l/N$. For simplicity, in order to select the acceptance thresholds of the PE tests [36], the average behaviour of a typical fiber and detector model is considered (see Supplemental Material II for further details). We assume an overall system loss of 30 dB, a dark count probability of $p_{\rm d}=6\times{}10^{-7}$ [37], and a misalignment error rate of $e_{\rm mis}=5\times 10^{-3}$ [12]. As in the BBM92 protocol, the EC leakage is set to $\lambda_{\rm EC}=1.19N_{\rm Z}h(\theta^{\rm th})$, $\theta^{\rm th}$ denoting the correctable QBER of a key block [38]. Lastly, we set again the overall failure probability to $\epsilon_{\rm PE}=4\times 10^{-16}$, and assume a common value for each individual error term contributing to it [39]. Similarly, we take $\epsilon_{\rm cor}=\epsilon_{\rm PA}=\delta=10^{-8}$, where $\delta>0$ is an intrinsic variable of the security analysis (see Supplemental Material II). Overall, this leads to $\epsilon_{\rm sec}={}7\times{}10^{-8}$.

### v3

@@ Abstract
for arbitrary values of $N$, $n$, $\epsilon$ and $p^{\rm th}$. Naturally, an interplay exists between $\epsilon$ and $q^{\rm th}$: fulfilling Eq. (1) with a very small $\epsilon$ comes at the price of tolerating larger values of $q^{\rm th}$. In the context of QKD —in which $\hat{q}$ is often referred to as the phase-error rate (PHER)— this translates into a direct penalty on the provably secure key length. Notably, the above formulation of the problem shows up exactly in ideal implementations of the BB84 [4] or the BBM92 [7, 8] protocols, and we adhere to the latter in the main text for didactic purposes. The detailed protocol description and the secret key length formula can be found in Appendix I. On the other hand, the popular decoy-state BB84 protocol [9, 10, 11, 12], which is implemented in most of today’s commercial QKD setups, is addressed in Appendix II.
@@
Crucially, our relaxation stems from a rational approximation of the Kullback-Leibler divergence, which is based on a logarithmic inequality [24] carefully selected to keep the CIs tight and handy at the same time. Indeed, higher-order rational approximations of the logarithm lead to much more convoluted formulas for the divergence. The reader is referred to Appendix III for the relaxed Chernoff bound, and the corresponding CIs are stated in Proposition 1 and Corollary 1 below.
@@
A proof of this claim is given in Appendix III. Importantly, since the claim is solely derived from the additive Chernoff bound, it automatically extends to a HG random variable [17, 21]. The upper extension goes as follows.
@@
On another note, the reader is referred to Appendix II for the failure probability estimation of a decoy-state QKD protocol, which relies not only on Proposition 2 but also on Proposition 1.
@@
On another note, in Appendix IV we include optimal one-sided CIs for the average of independent Bernoulli parameters [30, 31], potentially replacing $\Gamma^{\pm}_{n,\epsilon}(\hat{p})$ (Proposition 1) in the failure probability estimation of decoy-state QKD schemes. These CIs are related to the well-known Clopper-Pearson CIs of the binomial distribution, which in fact have been used in the context of QKD security before (see e.g. [32, 33]).
QKD simulations. Next, we show the impact that the proposed tools have in the finite-key performance of QKD. Let us consider the ideal BBM92 protocol first. Consistently with [15, 8], we evaluate a simplified scenario in which Alice and Bob share a $2N$-partite quantum state as a protocol input. Then, they jointly select between two binary and complementary quantum measurements uniformly at random on a round-by-round basis, in so each of them extracting a raw data block of $N$ bits. These blocks are randomly splitted into a fixed-length test (sifted-key) string of $n$ ($N-n$) bits. That is to say, both the key and the test string are randomly sampled from the two bases. In the simulations, we set the tolerated quantum bit error rate (QBER) to $p^{\rm th}=4.55\%$ for ease of comparison with [15], where this choice is motivated by the results of the Micius experiment [34]. Also, an error correction (EC) leakage of $\lambda_{\rm EC}=1.19(N-n)h(p^{\rm th})$ bits is assumed, $h(x)$ denoting the binary entropy of $x$. The correctness parameter, $\epsilon_{\rm cor}$, and the privacy amplification error, $\epsilon_{\rm PA}$, are both set to $10^{-8}$ for illustration purposes, and the failure probability bound is set to $\epsilon_{\rm PE}=4\times{}10^{-16}$. Altogether, these settings lead to an overall secrecy parameter of $\epsilon_{\rm sec}\approx{}5\times{}10^{-8}$ (see Appendix I).
@@
Let us now assess the more practical decoy-state BB84 protocol. We consider a scheme with three common intensities per basis, $\mu$ (signal), $\nu$ (decoy) and $\omega$ (vacuum), and employ the $Z$ ($X$) basis for key extraction (PE) with a fixed-size data block of $N_{\rm Z}$ ($N_{\rm X}$) bits drawn at random from the corresponding data pool. Again, in the simulations we contemplate blocks up to $N=N_{\rm Z}+N_{\rm X}=10^{6}$ bits, and superpose an inset of the minimum block-size regime. Naturally, we set $N_{\rm Z}=q_{\rm Z}^{2}/(q_{\rm Z}^{2}+q_{\rm X}^{2})N$, where $q_{\rm Z}$ ($q_{\rm X}$) denotes the $Z$ ($X$) basis probability, common to both parties. Also, while we fix $\omega=10^{-4}$ to account for the finite extinction ratio of the intensity modulator [35], the settings $\mu$ and $\nu$, together with their respective probabilities, $p_{\mu}$ and $p_{\nu}$, and the test basis probability, $q_{\rm X}$, are tuned to optimize the key length per block, $l/N$. For simplicity, in order to select the acceptance thresholds of the PE tests [36], the average behaviour of a typical fiber and detector model is considered (see Appendix II for further details). We assume an overall system loss of 30 dB, a dark count probability of $p_{\rm d}=6\times{}10^{-7}$ [37], and a misalignment error rate of $e_{\rm mis}=5\times 10^{-3}$ [12]. As in the BBM92 protocol, the EC leakage is set to $\lambda_{\rm EC}=1.19N_{\rm Z}h(\theta^{\rm th})$, $\theta^{\rm th}$ denoting the correctable QBER of a key block [38]. Lastly, we set again the overall failure probability to $\epsilon_{\rm PE}=4\times 10^{-16}$, and assume a common value for each individual error term contributing to it [39]. Similarly, we take $\epsilon_{\rm cor}=\epsilon_{\rm PA}=\delta=10^{-8}$, where $\delta>0$ is an intrinsic variable of the security analysis (see Appendix II). Overall, this leads to $\epsilon_{\rm sec}={}7\times{}10^{-8}$.